OpenStudy (anonymous):

What is the total cost of a pair of jeans that are priced $50.00 if there is 7 1/2% sales tax? $53.75 $55.50 $57.50

OpenStudy (anonymous):

The jeans are $50, and with 7.5% sales tax ($3.75) it gives you a total of $53.75.

OpenStudy (anonymous):

Cost price = $ 50. and Tax @ 7.5 % Hence The .Total cost = 50 + 7.5% of 50 TC = 50 ( 1 + 7.5/100 ) TC = 50 ( 1.075 ) => TC = $ 53.75 .......... Answer Answer

OpenStudy (anonymous):

Just had this question and got it wrong because of @GirlByte's answer. The correct answer is 0.05